# BERT or FastText? A Comparative Analysis of Contextual as well as Non-Contextual Embeddings

### v1

@@ 1 Introduction
Word embedding is a way of representing words into dense vectors in a continuous space such that the vectors capture the semantic relationship between the words for the models to understand the context and meaning of the text. FastText, a context-independent method, basically captures the subword information, enabling it to learn rare words, misspelled words as well as out-of-vocabulary words. It is recognized in the NLP community for its efficient performance in tasks like text classification and sentiment analysis. Despite being relatively old, it still remains one of the most effective alternatives when performing tasks on large datasets across various languages.
@@
BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2018), a recently emerged word embedding method, which understands the meaning of a word based on its context in a sentence. This has led to improved performance for various tasks.
FastText is widely used for low-resource languages due to its subword-based approach. On the other hand, numerous variations of BERT like IndicBERT Kakwani et al. (2020), MuRIL Khanuja et al. (2021), AfriBERT Ralethe (2020), and mBERT Devlin et al. (2018) to name a few, are available for experiments with applications involving Low Resource languages. Recent studies have experimented with both FastText and BERT for various tasks.
@@
Experiments of D’Sa et al. (2020) demonstrated that BERT embeddings outperformed FastText for classifying English text into toxic and non-toxic. Findings of Ahmed et al. (2024) suggested that BERT embeddings outperformed those of FastText with an F1 score of 84% when evaluated for depressive post-detection in Bangla.
@@
This paper focuses on utilizing FastText and BERT for the Marathi language for the following tasks: Sentiment Classification, 2-Class and 4-Class Hate Speech Detection, and News Article classification for headlines, long paragraphs and long documents. We construct a comprehensive analysis of FastText embeddings, IndicFT Kakwani et al. (2020) and MahaFT Joshi (2022a) embeddings and BERT embeddings, muril-base-cased Khanuja et al. (2021) and marathi-bert-v2 Joshi (2022a). Additionally, we experimented on compressed as well as contextual and non-contextual variants of the BERT-based model embeddings. This analysis revealed that contextual BERT embeddings outperformed FastText for almost all the tasks evaluated.
Section 2 provides a concise review of prior research on FastText and BERT. Section 3, includes the datasets and model embeddings that are utilized for the experiments. Section 4 presents the methodology used. Section 5 presents the results and key insights drawn from the findings along with a comparative analysis of FastText embeddings and BERT. In section 6, we conclude our discussion.
@@
Muril | 76.58 | 65.77 | 81.76 | 85.95 | 91.61 | 81.36 | |
Muril (Compressed) | 75.16 | 63.25 | 81.44 | 82.72 | 90.39 | 79.00 |
@@
2 Literature Review
In conclusion, while BERT consistently outperforms other word embeddings in tasks like news classification, sentiment analysis, and hate speech detection in high-resource languages (HRLs) like English, its efficacy on LRLs remains relatively underexplored. Specifically, the comparison of BERT and FastText embeddings for NLP tasks in LRLs like news classification, sentiment analysis, and hate speech detection still warrants further investigation to assess BERT’s effectiveness across a wider range of languages and tasks.
@@ 4 Methodology
For each sentence, corresponding embeddings were generated. The creation of BERT embeddings was done by first tokenizing the text using the BERT tokenizer, along with padding and truncation. The tokenized input was then passed to the model and the output of the last hidden layer of BERT was taken, which was then averaged to get contextual embeddings for every sentence. Whereas for non-contextual embeddings, the output of the first embedding layer was used.
However, for FastText, which is a non-contextual embedding by default, the process was slightly different due to the lack of a predefined vocabulary. Unlike BERT, which uses a tokenizer that can handle various Marathi words, FastText requires
the creation of a custom vocabulary. To achieve this, the training and validation datasets were concatenated and passed through a text vectorizer, which generated vectors for every word in the dataset. The vectorizer returned the vocabulary as a list of words in decreasing order of their frequency. The FastText model was then loaded using the FastText library, and for each word in the vocabulary, a word vector was retrieved to construct the embedding matrix. For each sentence, the text was split into individual words, and the corresponding embeddings were retrieved from the embedding matrix. These embeddings were then averaged to produce the final sentence embeddings.
@@
Singular Value Decomposition (SVD) to select the most relevant features, extracting the top 300 components for all the combinations of contextual as well as non-contextual for MahaBERT as well as Muril. All embeddings were then passed to a multiple logistic regression(MLR) classifier for classification into target labels. To determine the randomness, K-fold validation with 5 folds for all tasks
was performed, with the results laid out in Table 2.
@@ 5 Results
In Table 1, in the case of contextual embeddings, a decreasing trend is observed with the order: MahaBERT, Muril, MahaFT, and IndicFT. Similar results have also been observed for non-contextual embeddings. However, for non-contextual embeddings, a deviation in the above trend is observed in MahaSent and LDC datasets, where MahaFT outperforms MahaBERT.
From Table 2, we can infer that MahaSent results are inconclusive due to the high variance observed. However, we can conclude that LDC does show significant deviation from the trend, due to the relatively low variance.
Analyzing the t-SNE embedding visualizations for different datasets, BERT-based embeddings (MahaBERT and Muril) tend to form more compact clusters compared to FastText-based embeddings. Although some class overlap exists, BERT embeddings show denser class formations.
Furthermore, when evaluating the effect of compression, non-contextual embeddings tend to perform better in their uncompressed form, while no consistent trend was observed for compressed contextual embeddings.
@@ 6 Conclusion
In our research, we analyzed the effectiveness of various BERT and FastText-based embeddings on three key NLP tasks for Marathi: news classification, hate speech classification, and sentiment classification. Additionally, we examined the spatial distributions of embeddings generated for each classification task. Our study shows that MahaBERT contextual embeddings generally outperform MahaFT, Muril, and IndicNLP FastText embeddings across these tasks. This performance advantage is especially clear for contextual BERT embeddings, which show a strong trend of superiority. In contrast, non-contextual BERT embeddings generally follow this pattern but exhibit some deviations depending on the task.

### v2

@@ 1 Introduction
Word embedding is a way of representing words into dense vectors in a continuous space such that the vectors capture the semantic relationship between the words for the models to understand the context and meaning of the text. FastText, a context-independent method, basically captures the subword information, enabling it to learn rare words, misspelled words as well as out-of-vocabulary words. It is recognized in the NLP community for its efficient performance in tasks like text classification and sentiment analysis. Despite being relatively old, it still remains one of the most effective alternatives when performing tasks on large datasets across various languages due to its subword-based approach.
@@
BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2018) word embeddings understand the meaning of a word based on its context in a sentence. This has led to improved performance for various tasks. Numerous variations of BERT like IndicBERT Kakwani et al. (2020), MuRIL Khanuja et al. (2021), AfriBERT Ralethe (2020), and mBERT Devlin et al. (2018) to name a few, are available for experiments.
@@
Recent studies have experimented with both FastText and BERT for various tasks. Experiments of D’Sa et al. (2020) demonstrated that BERT embeddings outperformed FastText for classifying English text into toxic and non-toxic. Findings of Ahmed et al. (2024) suggested that BERT embeddings outperformed those of FastText with an F1 score of 84% when evaluated for depressive post-detection in Bangla.
While BERT consistently outperforms other word embeddings in various tasks in high-resource languages (HRLs) like English, its efficacy on low-resource languages (LRLs) remains relatively underexplored. Specifically, the comparison of BERT and FastText embeddings for tasks in LRLs like news classification, sentiment analysis, and hate speech detection still warrants further investigation to assess BERT’s effectiveness across a wider range of languages and tasks. This gap is particularly pronounced when balancing model performance with computational efficiency, which becomes a critical factor in low-resource settings.
Due to limited work in low-resource languages, it is not well studied as to which one will perform better in LRL scenarios. While BERT contextual embeddings are shown to work better than non-contextual FastText, it has significant computational overhead as compared to the FastText model. In our work, we focus on investigating how well non-contextual BERT embeddings from the first embedding layer of the model perform in comparison to FastText. Further, since BERT embeddings are 768-dimensional while FastText embeddings are 300-dimensional, we perform Singular Value Decomposition (SVD)-based compression on BERT embeddings to examine the performance of compressed BERT embeddings.
@@
This paper focuses on utilizing FastText and BERT for the Marathi language for the following tasks: Sentiment Classification, 2-Class and 4-Class Hate Speech Detection, and News Article Classification for headlines, long paragraphs, and long documents. We construct a comprehensive analysis of FastText embeddings, IndicFT Kakwani et al. (2020) and MahaFT Joshi (2022a) embeddings, and BERT embeddings, including muril-base-cased Khanuja et al. (2021) and marathi-bert-v2 Joshi (2022a). Additionally, we experimented on compressed as well as contextual and non-contextual variants of the BERT-based model embeddings. This analysis revealed that contextual BERT embeddings consistently outperformed FastText across all evaluated tasks. Additionally, non-contextual BERT embeddings demonstrated superior performance to FastText in most tasks.
@@
Examining the trend between non-contextual BERT and FastText embeddings.
The paper is organised as follows: Section 2 provides a concise review of prior research on FastText and BERT. Section 3 includes the datasets and model embeddings that are utilized for the experiments. Section 4 presents the methodology used. Section 5 presents the results and key insights drawn from the findings along with a comparative analysis of FastText embeddings and BERT. In Section 6, we conclude our discussion.
2 Literature Review
We note that the reviewed literature underscores the consistent superiority of BERT embeddings across diverse NLP tasks and domains. However, most existing work predominantly focuses on high-resource languages, emphasizing the need for further exploration and adaptation of these techniques for low-resource languages to achieve advancements in NLP.
@@ 2 Literature Review
| Muril | 76.58 | 65.77 | 81.79 | 85.95 | 91.61 | 81.36 | |
| Muril (Compressed) | 75.16 | 63.25 | 81.44 | 82.72 | 90.39 | 79.00 | |
@@ 4 Methodology
For each sentence, corresponding embeddings were generated. The creation of BERT embeddings was done by first tokenizing the text using the BERT tokenizer, along with padding and truncation. The tokenized input was then passed to the model and the output of the last hidden layer of BERT was taken, which was then averaged to get contextual embeddings for every sentence. Whereas for non-contextual embeddings, the output of the first embedding layer was used. Refer Fig 1 for the embedding extraction workflow diagram.
However, for FastText, which is a non-contextual embedding by default, the process was slightly different due to the lack of a predefined vocabulary. Unlike BERT, which employs a tokenizer capable of processing entire Marathi sentences, FastText necessitates the creation of a custom vocabulary. To achieve this, the training and validation datasets were concatenated and passed through a text vectorizer, which generated vectors for every word in the dataset. The vectorizer returned the vocabulary as a list of words in decreasing order of their frequency. The FastText model was then loaded using the FastText library, and for each word in the vocabulary, a word vector was retrieved to construct the embedding matrix. For each sentence, the text was split into individual words, and the corresponding embeddings were retrieved from the embedding matrix. These embeddings were then averaged to produce the final sentence embeddings.
@@
Singular Value Decomposition (SVD) to select the most relevant features, extracting the top 300 components for all the combinations of contextual as well as non-contextual for MahaBERT as well as Muril. Refer Fig 2 for compression flow diagram. All embeddings were then passed to a multiple logistic regression(MLR) classifier for classification into target labels.
To determine the validity of the results obtained, 5-fold cross-validation was performed for all tasks, and the results are presented in Table 2.
@@ 5 Results
Table 1 presents the results for various embeddings, including MahaBERT, MuRIL, MahaFT, and IndicFT, across multiple datasets and tasks. It includes both contextual and non-contextual embeddings, as well as the compressed variants of MahaBERT and MuRIL.
In sections 5.1 and 5.2, we have considered the uncompressed versions of Muril and MahaBERT. Further, in section 5.3, we specifically show the effect of compression on Muril and MahaBERT. In Section 5.4, we identify patterns in the embedding visualizations shown in Figure 3.
5.1 Contextual vs FastText
From Table 1, we observe the following trend when comparing contextual embeddings with FastText embeddings: MahaBERT > MuRIL > MahaFT > IndicFT.
5.2 Non-Contextual vs FastText
The trend of comparing non-contextual embeddings with FastText typically follows this order: MahaBERT > MuRIL > MahaFT > IndicFT. However, there are exceptions for the MahaSent and LDC datasets.
For these two datasets, FastText tends to perform slightly better. However, the difference is minimal, so we refer to Table 2 to determine whether this deviation is significant or simply random noise. We observe a high variance in MahaSent, suggesting that its deviation from the usual trend when comparing non-contextual embeddings with FastText may be attributed to noise and is unlikely to be significant.
In contrast, the LDC dataset also deviates from the trend but exhibits relatively low variance. As a result, for the LDC dataset, the performance trend when comparing non-contextual embeddings with FastText is as follows: MahaFT > IndicFT > MahaBERT > MuRIL.
5.3 Effect of Compression
From Table 1, we can also infer that compression has a negative effect on non-contextual embeddings, as uncompressed ones tend to perform better. However, the impact of compression on contextual embeddings varies across datasets, making it difficult to draw a consistent conclusion.
5.4 Inference from TSNE
Analyzing the t-SNE embedding visualizations for LDC specifically, BERT-based embeddings (MahaBERT and Muril) tend to form more compact clusters compared to FastText-based embeddings. Although some class overlap exists, BERT embeddings show denser class formations.
@@ 6 Conclusion
In our research, we analyzed the effectiveness of various BERT and FastText-based embeddings on three key NLP tasks for Marathi: news classification, hate speech classification, and sentiment classification. Additionally, we examined the spatial distributions of embeddings generated for one of the classification task.
Our study shows that contextual MahaBERT embeddings generally outperform contextual Muril as well as FastText-based embeddings across the 3 tasks. Our results indicate that contextual embeddings yield better outcomes than non-contextual ones .
Regarding the effect of compression: compressing non-contextual embeddings has a negative impact on performance, whereas compressing contextual embeddings shows no definitive effect.
When comparing non-contextual BERT embeddings to FastText embeddings, the non-contextual BERT embeddings outperform the FastText embeddings in the majority of tasks.